# AudioLens: A Closer Look at Auditory Attribute Perception of Large Audio-Language Models

### v1

@@ I Introduction
Our contributions are: (1) the first study of internal information processing in LALMs; (2) revealing layer-wise information dynamics and their relation to recognition accuracy; (3) analyzing information flow across tokens to identify the information sources for attribute predictions; and (4) introducing a novel improvement method based on these findings. Our work advances understanding of LALMs’ internal mechanisms and suggests directions for future enhancement.
@@ IV-A Preliminaries: Logit Lens
Consider an LLM with $L$ layers, hidden dimension $d$, and vocabulary $V$ of size $|V|$. To examine the information at token position $i$ in layer $\ell$, let $\mathbf{h}^{\ell}_{i}\in\mathbb{R}^{d}$ denote the hidden representation at position $i$ and layer $\ell$. Logit Lens projects $\mathbf{h}^{\ell}_{i}$ onto the vocabulary space via the model’s unembedding matrix [^1] The unembedding matrix is the language model (LM) head that maps the final-layer hidden representations to logits over the vocabulary, which are then converted into a probability distribution for next-token prediction. $W_{U}$in $\mathbb{R}^{|V|\times d}$, producing a vector of logits. Applying softmax yields a probability distribution over the vocabulary:
@@
The effectiveness of the Logit Lens technique stems from the residual stream in transformer models, where each layer adds information into the stream and promotes the probability of concepts it encodes [34]. Prior studies have shown that these intermediate probability distributions encode rich and interpretable internal information, including factual knowledge, attributes, and relational cues about entities [25, 34, 65]. For example, if position $i$ corresponds to the last token in a description of an entity $e$ [^2] For entities that span multiple tokens, a common practice is to use the first token as a representative [25]., then the probability $p^{\ell}_{i}(e)$ assigned to $e$ can serve as a proxy for how much information about $e$ is recoverable at layer $\ell$ when processing the description [25]. Note that the distribution $\mathbf{p}^{L}_{i}$ at the final layer $L$ matches the next-token probability distribution at position $i$.
@@ V-A Dataset
We focus on four auditory attributes: speaker gender, spoken language, speaker emotion, and the animal producing the sound. The dataset contains triplets comprising an audio input, a textual prompt, and the corresponding attribute label. The audio samples and attribute labels are sourced from the SAKURA benchmark [20], which provides 500 samples per attribute. There are 2, 8, 5, and 9 distinct labels for gender, language, emotion, and animal, respectively.
@@
P2 (QA) $\mathtt{<USER>}$What is the gender of the speaker in the speech?$\mathtt{<ASST>}$ The speaker’s gender is $\mathtt{<USER>}$What is the language spoken in the speech? $\mathtt{<ASST>}$The speech’s spoken language is $\mathtt{<USER>}$What is the emotion of the speaker in the speech? $\mathtt{<ASST>}$ The speaker’s emotion is $\mathtt{<USER>}$What animal makes the sound? $\mathtt{<ASST>}$The sound file’s animal is
@@ V-B Investigated Models
We investigate three open-source LALMs: DeSTA2 [7], Qwen-Audio-Chat (Qwen)[5], and Qwen2-Audio-Instruct (Qwen2)[6]. These models are selected for their strong performance on the attribute recognition tracks of SAKURA [20], from which we source our dataset. Additionally, they perform competitively on other speech and audio benchmarks [16, 17], making them well-suited for our analyses. We implement Logit Lens on these models using the Patchscopes toolkit [62].
@@ VI-B RQ2: Information Evolution in Correct/Wrong Predictions
TABLE III: Pearson correlation and p-value between accuracies and critical layers for three LALMs. Significant p-values ($<$0.05) are bolded. “Excluding Gender” indicates correlations computed without gender track data.
@@ VI-D RQ4: Information Flow Across Token Positions
To quantify this, we mask auditory inputs during self-attention [^4] For DeSTA2, the speech transcriptions in the inputs are also masked. at the last token [^5] Masking applies only at the last token; other positions are unaffected., forcing the model to rely solely on hidden representations at preceding text token positions for attribute prediction (gray lines in Fig. 4). In most cases, we observe a notable drop in information scores and prediction accuracies, showing that information aggregated at the preceding text token positions alone is insufficient, and LALMs heavily rely on information directly obtained from auditory inputs when making predictions.
@@ VI-E RQ5: Demonstration of Applications
Fig. 5 shows accuracy on 400 testing samples across different $\lambda$ values. Our method, selecting $\bar{\ell}$ via layer-wise information scores, significantly outperforms the baseline over a wide range of $\lambda$, demonstrating its ability to identify layers containing meaningful information. We also observe that the choice of $\lambda$ is critical, as both excessively small and large values result in suboptimal performance. With a proper scaling factor $\lambda$, our method achieves a relative accuracy improvement of 16.3% over the original performance of DeSTA2 (i.e., when $\lambda=0$), without requiring any additional training.

### v2

@@ I Introduction
Our contributions are: (1) the first study of internal information processing in LALMs; (2) revealing layer-wise information dynamics and their relation to recognition accuracy; (3) analyzing information flow across tokens to identify the information sources for attribute predictions; and (4) introducing a novel improvement method based on these findings. Our work advances understanding of LALMs’ internal mechanisms and suggests directions for future enhancement. Code will be available at https://github.com/ckyang1124/AudioLens.
@@ IV-A Preliminaries: Logit Lens
Consider an LLM with $L$ layers, hidden dimension $d$, and vocabulary $V$ of size $|V|$. To examine the information at token position $i$ in layer $\ell$, let $\mathbf{h}^{\ell}_{i}\in\mathbb{R}^{d}$ denote the hidden representation at position $i$ and layer $\ell$. Logit Lens projects $\mathbf{h}^{\ell}_{i}$ onto the vocabulary space via the model’s unembedding matrix [^1] The unembedding matrix is the language model (LM) head that maps the final-layer hidden representations to logits over the vocabulary, which are then converted into a probability distribution for next-token prediction. $W_{U}$ $\in$ $\mathbb{R}^{|V|\times d}$, producing a vector of logits. Applying softmax yields a probability distribution over the vocabulary:
@@
The effectiveness of the Logit Lens technique stems from the residual stream in transformer models, where each layer adds information into the stream and promotes the probability of concepts it encodes [34]. Prior studies have shown that these intermediate probability distributions encode rich and interpretable internal information, including factual knowledge, attributes, and relational cues about entities [25, 34, 65]. For example, if position $i$ corresponds to the last token in a description of an entity ee [^2] For entities that span multiple tokens, a common practice is to use the first token as a representative [25]., then the probability $p^{\ell}_{i}(e)$ assigned to $e$ can serve as a proxy for how much information about $e$ is recoverable at layer $\ell$ when processing the description [25]. Note that the distribution $\mathbf{p}^{L}_{i}$ at the final layer $L$ matches the next-token probability distribution at position $i$.
@@ V-A Dataset
We focus on four auditory attributes: speaker gender, spoken language, speaker emotion, and the animal producing the sound. The dataset contains triplets comprising an audio input, a textual prompt, and the corresponding attribute label. The audio samples and attribute labels are sourced from the SAKURA [^3] https://github.com/ckyang1124/SAKURA benchmark [20], which provides 500 samples per attribute. There are 2, 8, 5, and 9 distinct labels for gender, language, emotion, and animal, respectively.
@@
[…]
@@ V-B Investigated Models
We investigate three open-source LALMs: DeSTA2444https://github.com/kehanlu/DeSTA2 [7], Qwen-Audio-Chat [^5] https://github.com/QwenLM/Qwen-Audio (Qwen) [5], and Qwen2-Audio-Instruct [^6] https://github.com/QwenLM/Qwen2-Audio (Qwen2) [6]. These models are selected for their strong performance on the attribute recognition tracks of SAKURA [20], from which we source our dataset. Additionally, they perform competitively on other speech and audio benchmarks [16, 17], making them well-suited for our analyses. We implement Logit Lens on these models based on the Patchscopes toolkit [^7] https://pair-code.github.io/interpretability/patchscopes/ [62].
@@ VI-B RQ2: Information Evolution in Correct/Wrong Predictions
TABLE III: Pearson correlation and p-value between accuracies and critical layers for three LALMs. Significant p-values (<<0.05) are bolded. “Excluding Gender” indicates correlations computed without gender track data.
@@ VI-D RQ4: Information Flow Across Token Positions
To quantify this, we mask auditory inputs during self-attention [^9] For DeSTA2, the inputs include speech transcriptions, which we treat as part of the auditory input. To ensure consistency with other models that do not use transcriptions, we mask them during processing. at the last token [^10] Masking applies only at the last token; other positions are unaffected., forcing the model to rely solely on hidden representations at preceding text token positions for attribute prediction (gray lines in Fig. 4). In most cases, we observe a notable drop in information scores and prediction accuracies, showing that information aggregated at the preceding text token positions alone is insufficient, and LALMs heavily rely on information directly obtained from auditory inputs when making predictions.
@@ VI-E RQ5: Demonstration of Applications
Fig. 5 shows accuracy on 400 testing samples across different $\lambda$ values. Our method, selecting $\bar{\ell}$ via layer-wise information scores, significantly outperforms the baseline over a wide range of $\lambda$, demonstrating its ability to identify layers containing meaningful information. We also observe that the choice of $\lambda$ is critical, as both excessively small and large values result in suboptimal performance. With a proper scaling factor $\lambda$ (i.e., when $\lambda=1.25$ in Fig. 5), our method achieves a relative accuracy improvement of 16.3% over the original performance of DeSTA2 (i.e., when $\lambda=0$), without requiring any additional training.
@@ VII Conclusion
VIII Limitations
We acknowledge several limitations of this work. Our analysis of LALMs employs the Logit Lens technique, which is common in existing literature, though we note that alternative variants of this method exist. Furthermore, as the first study to examine the auditory information processing of LALMs, our investigation primarily focuses on their ability to perceive and recognize fundamental auditory attributes. While these capabilities form the foundation for more advanced reasoning over auditory modalities, the underlying reasoning processes of LALMs are likely to involve distinct, multi-layered, and potentially more intricate patterns than those observed in foundational attribute perception. A comprehensive examination of such processes, particularly how LALMs integrate auditory cues with linguistic and world knowledge to perform higher-level inference, remains an open question. We leave this in-depth exploration for future work.
IX Acknowledgement
We thank the reviewers for their constructive feedback during the review process, which greatly contributed to improving this work. We also acknowledge the computational and storage support provided by the National Center for High-performance Computing (NCHC) of the National Applied Research Laboratories (NARLabs) in Taiwan.